# From Words to Numbers: Your Large Language Model Is Secretly A Capable Regressor When Given In-Context Examples

### v1

@@ C.1.1 Prompt
For the experiments performed with $100$ random seeds and $50$ random input-output tuples (i.e., $\mathbb{D}_{50}$), we simply skip the invalid generations. This is not problematic because it rarely occurs. For example, it has never occurred for the any of the experiments showed in Section 3 and Section 4. Moreover, given that we run each experiment with $100$ random seeds, skipping the invalid generations will still leave us with enough samples to estimate the statistics. For example, out of the results presented in Section 3 and 4, the largest number of invalid generations for a given dataset was 11, by Mistral Medium on Regression NI 1/2. Claude 3 Opus produced only 2 invalid generations, for Friedman #2, and GPT-4 did not produce any invalid generations. One exception is for the results shown in Appendix H, as we observed that Claude 2.0 and Claude 2.1 produced invalid generations very often. For example, Claude 2.1 generated invalid generations for the dataset Regression 1/1 in 71% of the cases. The reason for invalid generations was usually because Claude 2.1 refused to give “potentially misleading numerical estimates without proper context”. A second exception is for Striped Hyena Nous 7B (Appendix P). We add a cross ($\times$) to the cell in the rank heatmap for every model-dataset pair where the number of valid generations is under $20$.
@@
C.2 Traditional Supervised Models
@@
    Multi-Layer Perceptron: We used 6 variants of multi-layer preceptrons: 3 with different widths (MLP Wide 1, MLP Wide 2, MLP Wide 3) and 3 with different depths (MLP Deep 1, MLP Deep 2, MLP Deep 3).
@@
    KNN: We used multiple variants of KNN, where we vary the number of neighbors, the type of distance used, and the power parameter for the Minkowski metric; We distinguish between them with a v{index}.
@@
We used the sklearn implementation for each model. [^12] We used sklearn 1.4.1.post1. Similar to the LLM case, we do not tune any hyperparameters. We use the default hyperparameters available in sklearn. We remark that these supervised baselines are very strong, as (1) many of them are the results of algorithms specifically designed for regression (e.g., Splines), (2) all perform parameter updates, and (3) the default hyperparameters, as set in widely-used statistical packages, have been refined over time to offer a reliable and generally strong performance across a variety of scenarios.
C.3 Unsupervised Models
@@
    Last: Uses the most recent observation tuple $(x_{n},y_{n})$ for prediction, such that $y_{n+1}=y_{n}$.
@@
Appendix D Large Language Models Can Do Linear Regression (Expanded)
@@
Fourth, there are specific LLMs which always perform better than Gradient Boosting, such as Claude 3 Opus and GPT-4. We remark that DBRX and Code Llama 70B outperform Gradient Boosting in 4 out of 6 datasets.
@@
Appendix E Large Language Models Can Do Non-Linear Regression (Expanded)
@@
Appendix F Average Model Ranks
@@
Appendix G How Fast Do Large Language Models Adapt? (Expanded)
@@
Appendix H Claude Performance
@@ Appendix I Costs
Appendix J LLMs Justifying Their Prediction
Without adding the prefix instruction text (see prompt in Appendix C.1.1), which instructed the models to output only its best estimate, some LLMs (e.g., Claude 3 Opus) started to provide explanations, in an attempt to justify their prediction. [^15] We hypothesize that this is because their system prompt instructs them to provide explanations. Analyzing these “explanations” revealed that there is a discrepancy between their explanation and their prediction. For example, for a non-linear regression problem, Claude 3 Opus suggests to train a Linear Regression model and gives the code to do so. Then it elaborates on how to use it for inference and gives the output. However, manually running the code suggested by the model results in a very different output. We provide some examples in Figures 3132, , 33, and 34. We describe each one below.
@@
Appendix K Beyond Numerical Regression
@@
Appendix L Effects of Rounding
@@
Appendix M Is It Just Better KNN?
We can see from Appendix D and E that the performance of certain LLMs is almost always better than that of all three variants of KNN presented. For example, in the case of linear regression, Claude 3 Opus, Claude 3 Sonet and GPT-4 always perform better than the three variants of KNN we used. Similar, in the case of non-linear regression, only for Simple NN 1 and Transformer 1 is any of the KNN variants outperforming Claude 3 Opus. Moreover, from Table 8, which records the best curve fit over the cumulative regret, we can see that both variants of KNN perform worse than Claude 3 Opus or GPT-4.
@@
Appendix N Could It Be Just Contamination?
The very large datasets that are used to train contemporary large language models (LLMs) raise concerns about potential (accidental or not) contamination Sainz et al. (2023); Golchin & Surdeanu (2024). In our study, we have attempted to mitigate this as follows. First, we used many different random seeds. However, this does not nullify the risk that the LLM has seen similar data (e.g., data from Friedman #1, but with other random seeds). To mitigate this, we explored the performance of the models on regression functions of our own creation. This makes it unlikely that the model has seen data coming from the exact same function. For example, across all our newly introduced datasets (e.g., Original #1, Original #2, Original #3, Original #4, Original #5), Claude 3 Opus obtains the highest average rank of $6.4$. Second place is Clade 3 Sonnet, with an average rank of $6.8$, then Gradient Boosting with $8.4$. Furthermore, our empirical evidence of consistent high performance across a diverse array of LLMs.
@@
N.1 LLMs with known Training Data
@@
N.2 Performance When Knowing The Dataset Name
@@
We present the comparative results in Table 9, which shows the average absolute error under conditions of dataset awareness versus unawareness. Notably, the mean absolute errors ($\downarrow$) remain closely matched across all scenarios. To statistically substantiate these observations, we performed paired t-tests for each dataset comparison. Given the multiplicity of tests performed, it became imperative to apply an adjustment for multiple comparisons to our p-values Dunn (1961); Benjamini & Hochberg (1995). Following this adjustment, none of the p-values remained below the (typically used) $0.05$ threshold, suggesting that the knowledge of the dataset name does not significantly affect model performance. Prior to adjustment, in two cases the resulting p-value was under $0.05$: GPT-4 on Friedman #3 (p-value $0.045$) and Claude 3 Sonnet on Friedman #3 (p-value $0.018$). Note, however, that only in the case of GPT-4 was the performance corresponding to the Dataset Aware setting better. This analysis indicates that, within the bounds of statistical significance, there is no substantial evidence to suggest that the performance of the models is influenced by explicit knowledge of the dataset name, something which has been linked to contamination (Golchin & Surdeanu, 2024). [^20] We investigated why the performance of Claude 3 Sonnet degraded on Friedman #2 when given the dataset name. We found that it is (mostly) because of an outlier: the absolute difference between the model’s prediction and expected output is $>300$.
@@
Appendix O Does The Performance Plateau?
@@
Appendix P Beyond Transformer-Based LLMs
In our study, we initially focused on transformer-based large language models (LLMs). To broaden our scope, we explore the capabilities of non-transformer LLMs, including a RWKV-based 14B LLM (Peng et al., 2023) and with StripedHyena Poli et al. (2023a; b), a 7B LLM. The performance rankings for these models, along with the transformer-based Mistral 7B for comparison, are illustrated in the heatmap provided in Figure 48. We tried running Falcon 7B as well, but it produced invalid outputs for almost all examples and all datasets, therefore we skip it. StripedHyena also encountered difficulties, producing invalid outputs in certain scenarios, such as 98% invalid responses for the Friedman #2 dataset. Consequently, we omitted Friedman #2 and Original #2 from its evaluation. However, it is important to highlight that the other models evaluated did not exhibit these issues and were able to generate valid outputs consistently. We make the following observations.

### v3

@@ D.1.1 Prompt
For the experiments performed with $100$ random seeds and $50$ random input-output tuples (i.e., $\mathbb{D}_{50}$), we simply skip the invalid generations. This is not problematic because it rarely occurs. For example, it has never occurred for the any of the experiments showed in Section 3 and Section 4. Moreover, given that we run each experiment with $100$ random seeds, skipping the invalid generations will still leave us with enough samples to estimate the statistics. For example, out of the results presented in Section 3 and 4, the largest number of invalid generations for a given dataset was 11, by Mistral Medium on Regression NI 1/2. Claude 3 Opus produced only 2 invalid generations, for Friedman #2, and GPT-4 did not produce any invalid generations. One exception is for the results shown in Appendix J, as we observed that Claude 2.0 and Claude 2.1 produced invalid generations very often. For example, Claude 2.1 generated invalid generations for the dataset Regression 1/1 in 71% of the cases. The reason for invalid generations was usually because Claude 2.1 refused to give “potentially misleading numerical estimates without proper context”. A second exception is for Striped Hyena Nous 7B (Appendix R). We add a cross ($\times$) to the cell in the rank heatmap for every model-dataset pair where the number of valid generations is under $20$.
@@
D.2 Traditional Supervised Models
@@
    Multi-Layer Perceptron: We used 6 variants of multi-layer preceptrons: 3 with different widths (MLP Wide 1, MLP Wide 2, MLP Wide 3) and 3 with different depths (MLP Deep 1, MLP Deep 2, MLP Deep 3)
@@
    KNN: We used multiple variants of KNN, where we vary the number of neighbors, the type of distance used, and the power parameter for the Minkowski metric; We distinguish between them with a v{index}
@@
We used the sklearn implementation for each model. [^12] We used sklearn 1.4.1.post1. Similar to the LLM case, we do not tune any hyperparameters. We use the default hyperparameters available in sklearn. We remark that these supervised baselines are very strong, as (1) many of them are the results of algorithms specifically designed for regression (e.g., Spline), (2) all perform parameter updates, and (3) the default hyperparameters, as set in widely-used statistical packages, have been refined over time to offer a reliable and generally strong performance across a variety of scenarios.
D.3 Unsupervised Models
@@
    Average: Predicts the next value, $y_{n+1}$, as the mean of all preceding outcomes: $y_{n+1}=\frac{1}{n}\sum_{i=1}^{n}y_{i}$
@@
    Last: Uses the most recent observation tuple $(x_{n},y_{n})$ for prediction, such that $y_{n+1}=y_{n}$
@@
Appendix E Large Language Models Can Do Linear Regression (Expanded)
@@
Fourth, there are specific LLMs which always perform better than Gradient Boosting, such as Claude 3 Opus and GPT-4. We remark that DBRX and Code Llama 70B outperform Gradient Boosting in 4 out of 6 datasets.  [^13] In order to see how the performance of LLMs scale with more variables, we additionally experimented with Claude 3 Opus on Regression NI 1/10, 3/10, and 5/10 respectively. For example, in NI 5/10, Claude 3 Opus outperforms methods like Gradient Boosting or Random Forest but lags behind methods like Logistic Regression, which is expected given the nature of the datasets. For NI 3/10, Claude 3 Opus is outperformed by Random Forest and Gradient Boosting. Overall, we observed the performance to be similar to the performance on NI 2/3 from Figure 16.
@@
Appendix F Large Language Models Can Do Non-Linear Regression (Expanded)
@@
Appendix G Average Model Ranks
@@
Appendix H How Fast Do Large Language Models Adapt? (Expanded)
@@
Appendix I Performance on Real-World Datasets
In the following, we include the performance of GPT-4 and Claude 3 Opus on real-world datasets. Nevertheless, our aim in this work has not been to suggest that LLMs should replace (at least currently) traditional regression methods. [^14] Although we can foresee some interesting advantages for LLMs, such as allowing for natural language description of what each feature represents. We evaluate on the following datasets:
  1. 1.
    Liver Disorders (UCI id=$60$ (UCI, 1990)
  2. 2.
    Real Estate Valuation (UCI id=$477$ (Yeh, 2018))
  3. 3.
  4. 4.
    Servo (UCI id=$87$ (Ulrich, 1993))
  5. 5.
    Movies (UCI id=$424$ (Ahmed, 2017))
We evaluated the following models:
  • LLMs: GPT-4, Claude 3 Opus, Claude 3 Sonnet, Claude 3 Haiku;
  • Traditional Supervised Methods: AdaBoost, Bagging, Gradient Boosting, KNN, KNN v2, KNN v3, Kernel Ridge, Lasso, Linear Regression, Linear Regression + Poly, MLP Deep 1, MLP Deep 2, MLP Deep 3, MLP Wide 1, MLP Wide 2, MLP Wide 3, Random Forest, Ridge, SVM, SVM + Scaler, Spline;
  • Unsupervised Methods: Average, Last, Random.
We show our results in Tables 9101112, and 13. In order to keep the tables easy to read, we selected: (i) two LLMs, Claude 3 Opus and GPT-4; (ii) three traditional supervised methods, Gradient Boosting, Random Forest, and Linear Regression with Polynomial Features (LR + Poly); and (iii) three unsupervised methods, Average, Last, and Random. We show both the mean absolute error (MAE) and the rank. The rank is shown against all the methods ran. For example, on Liver Disorders (Table 9), SVM ranks first.
Model MAE Rank
GPT-4 2.55$\pm$2.49 6
Gradient Boosting 2.57$\pm$1.95 8
Random Forest 2.62$\pm$1.57 12
Linear Regression + Poly 2.67$\pm$1.61 19
Average 2.79$\pm$1.60 22
Claude 3 Opus 3.04$\pm$2.13 24
Random 3.23$\pm$2.71 26
Last 3.46$\pm$2.82 27
Table 9: Results on the Liver Disorders dataset, sorted by the MAE ($\downarrow$)
Model MAE Rank
Gradient Boosting 4.30$\pm$4.91 1
Claude 3 Opus 5.08$\pm$5.35 4
GPT-4 5.25$\pm$5.88 6
Linear Regression + Poly 5.38$\pm$5.78 7
Random Forest 5.72$\pm$5.32 10
Average 11.12$\pm$7.69 26
Random 14.39$\pm$10.05 27
Last 15.27$\pm$11.67 28
Table 10: Results on the Real Estate Valuation dataset, sorted by the MAE ($\downarrow$)
Model MAE Rank
Random Forest 41.31$\pm$27.68 1
Gradient Boosting 47.95$\pm$28.36 16
GPT-4 50.26$\pm$45.25 17
Claude 3 Opus 54.13$\pm$39.71 19
Linear Regression + Poly 55.84$\pm$52.87 20
Average 68.02$\pm$37.57 24
Random 94.34$\pm$63.17 26
Last 103.96$\pm$69.57 27
Table 11: Results on the Diabetes dataset, sorted by the MAE ($\downarrow$)
Model MAE Rank
Gradient Boosting 0.25$\pm$0.29 2
Claude 3 Opus 0.27$\pm$0.46 4
GPT-4 0.44$\pm$0.62 10
Random Forest 0.45$\pm$0.44 11
Linear Regression + Poly 0.62$\pm$0.38 13
Average 1.43$\pm$1.27 24
Last 1.61$\pm$1.56 26
Random 1.77$\pm$1.74 27
Table 12: Results on the Servo dataset, sorted by the MAE ($\downarrow$)
Model MAE (1e+07) Rank
Random Forest 3.02$\pm$2.48 2
Gradient Boosting 3.03$\pm$2.51 3
Claude 3 Opus 4.51$\pm$6.20 17
GPT-4 5.48$\pm$7.21 20
Average 6.20$\pm$4.85 21
Linear Regression + Poly 6.40$\pm$7.10 22
Random 8.17$\pm$10.9 23
Last 8.26$\pm$6.22 24
Table 13: Results on the Movies dataset (scaled by 1e+07), sorted by the MAE ($\downarrow$)
Appendix J Claude Performance
@@ Appendix I Performance on Real-World Datasets
Appendix L LLMs Justifying Their Prediction
Without adding the prefix instruction text (see prompt in Appendix D.1.1), which instructed the models to output only its best estimate, some LLMs (e.g., Claude 3 Opus) started to provide explanations, in an attempt to justify their prediction. [^18] We hypothesize that this is because their system prompt instructs them to provide explanations. Analyzing these “explanations” revealed that there is a discrepancy between their explanation and their prediction. For example, for a non-linear regression problem, Claude 3 Opus suggests to train a Linear Regression model and gives the code to do so. Then it elaborates on how to use it for inference and gives the output. However, manually running the code suggested by the model results in a very different output. We provide some examples in Figures 3132, , 33, and 34. We describe each one below.
@@
Appendix M Beyond Numerical Regression
@@
Appendix N Effects of Rounding
@@
Appendix O Is It Just Better KNN?
We can see from Appendix E and F that the performance of certain LLMs is almost always better than that of all three variants of KNN presented. For example, in the case of linear regression, Claude 3 Opus, Claude 3 Sonet and GPT-4 always perform better than the three variants of KNN we used. Similar, in the case of non-linear regression, only for Simple NN 1 and Transformer 1 is any of the KNN variants outperforming Claude 3 Opus. Moreover, from Table 8, which records the best curve fit over the cumulative regret, we can see that both variants of KNN perform worse than Claude 3 Opus or GPT-4.
@@
Appendix P Could It Be Just Contamination?
The very large datasets that are used to train contemporary large language models (LLMs) raise concerns about potential contamination (Sainz et al., 2023; Golchin & Surdeanu, 2024). In our study, we have attempted to mitigate this as follows. First, we used many different random seeds. However, this does not nullify the risk that the LLM has seen similar data (e.g., data from Friedman #1, but with other random seeds). To mitigate this, we explored the performance of the models on regression functions of our own creation. This makes it unlikely that the model has seen data coming from the exact same function. For example, across all our newly introduced datasets (e.g., Original #1, Original #2, Original #3, Original #4, Original #5), Claude 3 Opus obtains the highest average rank of $6.4$. Second place is Clade 3 Sonnet, with an average rank of $6.8$, then Gradient Boosting with $8.4$. Furthermore, our empirical evidence of consistent high performance across a diverse array of LLMs.
@@
P.1 LLMs With Known Training Data
@@
P.2 Performance When Knowing The Dataset Name
@@
We present the comparative results in Table 14, which shows the average absolute error under conditions of dataset awareness versus unawareness. Notably, the mean absolute errors ($\downarrow$) remain closely matched across all scenarios. To statistically substantiate these observations, we performed paired t-tests for each dataset comparison. Given the multiplicity of tests performed, it became imperative to apply an adjustment for multiple comparisons to our p-values (Dunn, 1961; Benjamini & Hochberg, 1995). Following this adjustment, none of the p-values remained below the (typically used) $0.05$ threshold, suggesting that the knowledge of the dataset name does not significantly affect model performance. Prior to adjustment, in two cases the resulting p-value was under $0.05$: GPT-4 on Friedman #3 (p-value $0.045$) and Claude 3 Sonnet on Friedman #3 (p-value $0.018$). Note, however, that only in the case of GPT-4 was the performance corresponding to the Dataset Aware setting better. This analysis indicates that, within the bounds of statistical significance, there is no substantial evidence to suggest that the performance of the models is influenced by explicit knowledge of the dataset name, something which has been linked to contamination (Golchin & Surdeanu, 2024). [^23] We investigated why the performance of Claude 3 Sonnet degraded on Friedman #2 when given the dataset name. We found that it is (mostly) because of an outlier: the absolute difference between the model’s prediction and expected output is $>300$.
@@
Appendix Q Does The Performance Plateau?
@@
Appendix R Beyond Transformer-Based LLMs
In our study, we initially focused on transformer-based large language models (LLMs). To broaden our scope, we explore the capabilities of non-transformer LLMs, including a RWKV-based 14B LLM (Peng et al., 2023) and with StripedHyena (Poli et al., 2023a; b), a 7B LLM. The performance rankings for these models, along with the transformer-based Mistral 7B for comparison, are illustrated in the heatmap provided in Figure 48. We tried running Falcon 7B as well, but it produced invalid outputs for almost all examples and all datasets, therefore we skip it. StripedHyena also encountered difficulties, producing invalid outputs in certain scenarios, such as 98% invalid responses for the Friedman #2 dataset. Consequently, we omitted Friedman #2 and Original #2 from its evaluation. However, it is important to highlight that the other models evaluated did not exhibit these issues and were able to generate valid outputs consistently. We make the following observations.